A sample of nitrogen gas occupies a volume of 1.6 L at 838 mm Hg and -3° C. The volume increases by 1.3 L and the temperature decreases to 54 K. What is the final pressure exerted on the gas in mm Hg? Round to two decimal places.
1
Expert's answer
2019-03-15T02:35:25-0400

P1V1/T1 = P2V2/T2

P2 = P1V1T2/V2T1

-3° C = 270.15°K

P2 = (838*1.6*54)/(1.3*270.15) = 206.16 (mm Hg)